### v1

# BINDy – Bayesian identification of nonlinear dynamics with reversible-jump Markov-chain Monte-Carlo

Abstract
Model parsimony is an important cognitive bias in data-driven modelling that aids interpretability and helps to prevent over-fitting. Sparse identification of nonlinear dynamics (SINDy) methods are able to learn sparse representations of complex dynamics directly from data, given a basis of library functions. In this work, a novel Bayesian treatment of dictionary learning system identification, as an alternative to SINDy, is envisaged. The proposed method – Bayesian identification of nonlinear dynamics (BINDy) – is distinct from previous approaches in that it targets the full joint posterior distribution over both the terms in the library and their parameterisation in the model. This formulation confers the advantage that an arbitrary prior may be placed over the model structure to produce models that are sparse in the model space rather than in parameter space. Because this posterior is defined over parameter vectors that can change in dimension, the inference cannot be performed by standard techniques. Instead, a Gibbs sampler based on reversible-jump Markov-chain Monte-Carlo is proposed. BINDy is shown to compare favourably to ensemble SINDy in three benchmark case-studies. In particular, it is seen that the proposed method is better able to assign high probability to correct model terms.
Keywords Bayesian Statistics  $\cdot$ Nonlinear dynamics  $\cdot$ Reversible-jump  $\cdot$ MCMC
1 Introduction
Describing the dynamics of nonlinear processes analytically is of fundamental interest to many branches of scientific modelling. In cases where governing differential equations are unavailable, the practitioner has little choice but to try and discern differential equations directly from data. This difficulty is compounded in situations whereby both the model structure (which terms should be included) and its parameterisation are unknown. In high-noise and low-data regimes, model uncertainty quantification (UQ) becomes critical. Limited data, measurement noise, and an unknown model structure all contribute significant uncertainty.
An important facet of data-driven modelling is model parsimony. Recently, there has been increasing interest in methods that are able to develop parsimonious models directly from measured data. Methods such as sparse identification of nonlinear dynamics (SINDy) [1] have been extremely effective at learning sparse representations of complex dynamics. However, a key feature of all SINDy approaches is the specification of heuristic sparsity-inducing methods and hyperparameters that control the inevitable trade-off between fidelity and parsimony in data-driven modelling.
In this paper, the authors take the view that developing parsimonious models from data is inherently an uncertain task. Parsimonious model selection in practice requires interpretable answers to questions in the vein of ‘what is the probability that a given term should be included given the data?’ and ‘if one were to include this term, how is its value distributed?’. In order to access principled answers to the above, authors propose an alternative to SINDy within a Bayesian framework.
Bayesian uncertainty quantification is a natural framework for conducting UQ in data-driven modelling. Indeed, much has been written on the topic, e.g. [2, 3]. Physical knowledge and inductive biases (such as parsimony) can be incorporated as prior knowledge, making such modelling assumptions explicit. Indeed, many literature contributions propose Bayesian methods for UQ in both model parameterisation [4, 5, 6] and model structure [7, 8, 9, 10]. Further discussion of specific related works is presended in Section 3.4.
1.1 Contribution
In this work, interpretable prior distributions are placed over terms in the model (independently of their parameterisation) and over parameters (independently of their inclusion). This prior structure invokes a joint posterior distribution that can be marginalised to obtain useful distributions, that can robustly address modelling questions such as those introduced above.
The key contribution of this work is a novel Bayesian approach to the identification of nonlinear dynamics from a library of basis terms. In particular,
  • Heuristic sparsity-inducing regression is replaced with interpretable prior distributions over models – all modelling assumptions are made up-front.
  • An efficient sampler is proposed to produce samples from the target joint posterior distribution over model terms and parameters.
  • The proposed approach is shown to compare favourably to ensemble-SINDy in three case studies including a popular population dynamics dataset consisting of only 21 data points.
The remainder of this paper is structured as follows: The following section summarises the SINDy method and holds some discussion on parsimony in data-driven modelling. A third section reviews the reversible-jump theory and introduces the proposed approach and its position in the literature. A fourth section compares the proposed approach to the ensemble SINDy method on two benchmark case studies. A final section presents some discussion and directions for future work.
2 SINDy methods
The SINDy method proposed in 2016 by Brunton et al. [1] has received considerable attention in the scientific literature as a computationally-efficient way to learn differential equations directly from data. Since it was first proposed, the method has been proven to be effective in correctly identifying governing equations from both simulated and real-world data.
At their core, SINDy methods make the key assumption that the observed dynamics admit a first-order formulation that is linear in the parameters of a library of user-specified basis functions $\Theta$. Let $f$ denote the first-order ordinary differential equation (ODE) that describes the evolution of some state-space dynamics $x(t)$, [^1] The method can trivially be extended to systems of ODEs, PDEs etc. but for notational simplicity, the exposition in this paper will consider scalar $x$ and $f$ only.
$\dot{x}(t)=f(x(t))$ (1)
where the overdot represents a time derivative. Now the foundational assumption of the SINDy method can be stated,
$\dot{x}\approx\Theta(x)\Xi$ (2)
where $x$ is the state vector, $\Theta$ is a library of user-selected basis functions computed from $x$. Critically, the coefficient vector $\Xi$ is assumed to be sparse, whereby many elements of $\Xi$ are equal to 0. In practice, it is not often the case that measurements of both $x$ and $\dot{x}$ are available. It is usual in the SINDy literature to assume that only the states $x$ are measured and their derivatives $\dot{x}$ can be computed numerically (with sufficient accuracy such that (2) holds). The SINDy framework (for identification of ODEs) can thus be summarised by three steps:
  1. 1.
    The selection of a library of candidate basis terms.
  2. 2.
    The selection of a numerical differentiation scheme to compute $\dot{x}$.
  3. 3.
    Sparse regression to the coefficient vector $\Xi$ by heuristic means.
It is clear that a great deal of methods can be applied to each of the steps above. In this way, SINDy-type methods have come to encompass a family of approaches. Since the original SINDy algorithm was introduced in [1], extensions have been proposed to partial differential equation (PDE) systems [11], implicit dynamics [12], discrete dynamics and weak PDE solutions [13]. Furthermore, many methods have been applied to the sparse regression task including sequentially-thresholded least-squares [1], sparse relaxed regression [14] and forward orthogonal least-squares regression (FROLS) [15]. Several approaches also include uncertainty quantification such as the ensemble formulation in [16] and sparsity-inducing Bayesian methods [17, 10, 18]. Many of these methods have been made readily available to practitioners via an open-source python library [19].
3 Proposed approach
Of the three components in SINDy methods described in the previous section, this paper is concerned only with the third—determination of appropriate model terms and their parameters. At this stage, it is useful to present some discussion as to the role of parsimony in data-driven modelling.
3.1 Parsimony in data-driven modelling
Inherently, model parsimony is a cognitive bias, injected by modellers to promote interpretability and to prevent overfitting. In many sparsity-promoting methods (such as SINDy), parsimony is enforced by proxy; the importance of each term in the library is related to the corresponding size of the parameter in $\Xi$. Although this is a convenient proxy for parsimony, it is not without limitations. A critical limitation in practice is that small parameters in $\Xi$ do not necessarily correspond to small effects in the dynamics. Consider the effect of neglecting a small but negative damping term in the model leading to instability in extrapolation. Practitioners cannot know in advance how small parameters will behave in regions of the state that are not observed in the training data. Normalisation of the columns of $\Theta$ can help to address this shortcoming, but can also introduce ambiguity between correlated library functions. For a motivating example consider the problem of selecting between the terms $x^{2}$ and $x^{4}$, $x\in[0,1]$ in the presence of noise. If all terms in $\Theta(x)$ normalised to unit standard deviation, the difference between the two terms might appear trivial in the regime of the training data; the difference in extrapolation is evident. Introducing correlations of this form can ultimately harm model parsimony.
A particular issue with parameter size as a proxy for term importance is the selection of a threshold value, below which model terms should be excluded. In practice, this level can be difficult to select in a principled manner, leading to a spectrum of models at different levels of sparsity.
Alternative proxies for term inclusion have been considered based on greedy reduction of some error metric. A key example is forward regression least-squares optimisation (FROLS) [15, 19] that attempts to greedily add (or remove) model terms that maximize an error-reduction ratio. Such approaches circumvent problems of small parameter values described above. In practice however, these methods require the specification of a convergence threshold (or directly number the number of terms in the model, as in the implementation in [19]) which can also be difficult to set in advance in a principled manner.
It is the position of the authors that data alone cannot inform model parsimony. In order to select models that are parsimonious, cognitive biases must be applied. In this work, a wider probabilistic view of parsimony in SINDy-type methods is taken. Rather than use parameter magnitude as a proxy for sparsity, the authors propose to target the parameter inclusion probability directly. Taking a Bayesian view of (2) and assuming an additive Gaussian noise model one has,
$\dot{x}=\Theta(x)\Xi+\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\sigma^{2})$ (3)
where $\sigma^{2}$ is the noise variance in the observation model. The model identification task at hand is to identify a subset of the columns of $\Theta(x)$. Intuitively, this model set includes all subsets of library functions (the powerset) and thus all possible ways that the parameter vector may be sparsified. Let,
=$\mathcal{M}=\{m_{i}\}_{i=1}^{2^{n}}$ (4)
be the set of all such possible models for $\Theta(x)$ consisting of $n$ library functions [^2] The method exposed in this work does not require $n$ to be finite and can be generalised to the case of infinite libraries (for example all polynomials), however the exposition and results herein consider finite $n$ only. Some additional discussion is held in Section 5.. Key to the formulation here, is that each model in $m\in\mathcal{M}$ is parametrised by a corresponding parameter vector $\Xi$, the dimension of which depends on the model $m$.
The objective of this work is to infer the joint posterior distribution $p(\Xi,m|\dot{x},\Theta(x))$. The observation model in (3) gives rise to a Gaussian likelihood $p(\Xi,m|\dot{x},\Theta(x))$ for each model $m$ and corresponding parameter vector $\Xi$. With an appropriately defined prior $p(\Xi,m)$, and applying Bayes rule, one may write,
$p(\Xi,m|\dot{x},\Theta(x))=\frac{p(\Xi,m)p(\dot{x}|\Xi,m,\Theta(x))}{p(\dot{x})}$ (5)
For notational simplicity in the exposition that follows, the dependence of the likelihood and posterior on $\Theta(x)$ will be dropped, and the observed (or computed) state derivatives are denoted by $\mathcal{D}=\dot{x}$. The above can thus be written,
$p(\Xi,m|\dot{x},\mathcal{D})=\frac{p(\Xi,m)p(\mathcal{D}|\Xi,m)}{p(\mathcal{D})}$ (6)
Access to the posterior distribution over both the model terms and its parameterisation confers a number of advantages for the data-driven discovery of nonlinear dynamics. If samples from the posterior are available, a major advantage is that they can be marginalised (in the Monte-Carlo sense) to give direct access to a posterior distribution over models, independently of their parameterisation i.e. $p(m|\mathcal{D})$. This allows the practitioner to evaluate the probability that a term should be included in a model (given the observed data). This permits more robust ways to introduce model parsimony, for example by excluding terms with an inclusion probability below a certain threshold. Alternatively, the full posterior over model terms and parameters can be propogated to further analyses.
The joint posterior distribution $p(\Xi,m|\dot{x},\mathcal{D})$ over both models and their parameters is a very challenging object to approach. As is usual in complex Bayesian inference tasks, the model evidence term $p(\mathcal{D})$ is unavailable analytically. However, this is not the only obstacle. The major inferential challenge is that the dimension of the parameter vector $\Xi$ has come to depend on the particular model it parameterises. In order to overcome these difficulties, a sampling scheme based on reversible-jump Markov-chain Monte-Carlo (RJMCMC) [20] will be employed.
3.2 Reversible-jump MCMC
It will be useful here to briefly review both Metropolis-Hastings (MH) and reversible-jump Markov-chain Monte-Carlo theory in a general setting in order to motivate the proposed approach. Much of the exposition here is available in additional detail in references [21] for MH and [20] for reversible-jump MCMC. Consider the well-studied problem of sampling from a target distribution $\pi(\theta)$ (that is known up to some constant), by constructing a Markov chain with $\pi$ as its stationary distribution. A sufficient condition for convergence is that of detailed balance, whereby,
$\pi(\theta^{\prime})\kappa(\theta|\theta^{\prime})=\pi(\theta)\kappa(\theta^{\prime}|\theta)$ (7)
Where $\kappa(\theta^{\prime}|\theta)$ is the kernel of the Markov chain moving from state $\theta$ to new location $\theta^{\prime}$. The MH algorithm further divides this kernel into a transition density $p(\theta^{\prime}|\theta)$ and an acceptance probability $\alpha(\theta\rightarrow\theta^{\prime})$,
$k(\theta^{\prime}|\theta)=p(\theta^{\prime}|\theta)\alpha(\theta\rightarrow\theta^{\prime})$ (8)
The familiar MH acceptance probability can be found by maximising the acceptance probability while retaining the condition of detailed balance. The optimal choice is found to be,
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta)}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{p(\theta|\theta^{\prime})}{p(\theta^{\prime}|\theta)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}$ (9)
whereby an unknown normalisation constant in $\pi$ can be cancelled from the numerator and denominator.
In the case that $\theta^{\prime}$ and $\theta$ have different dimensions, detailed balance will not hold for standard choices of the transition density. To overcome this issue, in 1995 Green introduced the reversible-jump MCMC (RJMCMC) algorithm [20] as a method to sample from distributions defined over parameters of different dimension. The RJMCMC method overcomes this difficulty via dimension matching. Let $k$ be the dimension of the state of a Markov chain at $\theta$. In order to move to a new state with dimension $k^{\prime}$ and position $\theta^{\prime}$, an auxiliary variable $u^{\prime}$ with dimension $j^{\prime}$ is sampled. To compensate for the mismatch between dimensions, it is required that $j+k=j^{\prime}+k^{\prime}$, where $j$ is the current dimension of the auxillary variable $u$. Then, a bijective map $g_{k\rightarrow k^{\prime}}:\{u,\theta\}\rightarrow\{u^{\prime},\theta^{\prime}\}$ between each pair of dimensions is defined, such that the dimension of $\{u,\theta\}$ is unchanged by the bijection. If the probability of ‘jumping’ from one dimension to another (via the appropriate corresponding bijection) is given by $J(k^{\prime},\theta^{\prime}|k,\theta)q(u^{\prime}|\theta^{\prime})$ Then the RJMCMC acceptance probability can be reformulated as,
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta^{\prime})}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(k^{\prime},\theta^{\prime}|k,\theta)}{J(k,\theta|k^{\prime},\theta^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(u^{\prime}|\theta^{\prime})}{q(u|\theta)}}_{\begin{subarray}{c}\text{Auxillary}\\ \text{ratio}\end{subarray}}\underbrace{\left|\frac{\partial g_{k\rightarrow k^{\prime}}}{\partial(u,\theta)}\right|}_{\begin{subarray}{c}\text{Jacobian}\\ \text{determinant}\end{subarray}}\Big{\}}$ (10)
In the general setting, working with RJMCMC can be cumbersome for the practitioner. The specification of the bijective maps is non-trivial and has a strong effect on the sampling efficiency of the scheme. Fortunately, a significant simplification is sometimes available and will be employed here. The trick is to sample model parameters independently between models. This is achieved by stetting $u\triangleq\theta^{\prime}$ and $u^{\prime}\triangleq\theta$ in each transition. Thus, each bijection can be defined,
$g_{k\rightarrow k^{\prime}}(u,\theta)=\{u^{\prime},\theta^{\prime}\}=\{\theta,\theta^{\prime}\}$ (11)
Interrogating the above, it is clear that the dimension matching constraint is satisfied and that the Jacobian is a row-wise re-ordering of the identity matrix. Therefore, the Jacobian determinant term in the acceptance ratio must be identically equal to one. The acceptance probability may now be simplified,
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta^{\prime})}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(k^{\prime},\theta^{\prime}|k,\theta)}{J(k,\theta|k^{\prime},\theta^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(\theta|\theta^{\prime})}{q(\theta^{\prime}|\theta)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}$ (12)
3.3 Bayesian identification of nonlinear dynamics
With the relevant RJMCMC theory established, attention can be returned to the problem of sampling from the posterior density $p(\Xi,m|\mathcal{D})$. In order to develop the posterior over both the model terms and their parameterisation prior and proposal densities must be specified. It is clear that there are many appropriate choices for these objects and in practice it can be expected that prior selection will be guided by the problem at hand. However, it can be shown that certain choices of the parameter proposal can lead to drastic simplification of the inference scheme.
Considering for now the case whereby the noise variance $\sigma^{2}$ in (3) is known and substituting the required posterior distribution into the acceptance probability above, one has,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(\Xi^{\prime},m^{\prime}|\mathcal{D},\sigma^{2})}{p(\Xi,m|\mathcal{D},\sigma^{2})}}_{\begin{subarray}{c}\text{Posterior}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(\Xi|\Xi^{\prime})}{q(\Xi^{\prime}|\Xi)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}$ (13)
The proposal density in the above is freely chosen by the user. A natural choice, (after Troughton and Godsill [22]) is to employ the full conditional density,
$q(\Xi^{\prime}|\Xi)=p(\Xi^{\prime}|m^{\prime},\mathcal{D},\sigma^{2})$ (14)
substituting into the above, one finds that since,
$p(m|\mathcal{D},\sigma^{2})=\frac{p(\Xi,m|\mathcal{D},\sigma^{2})}{p(\Xi|m,\mathcal{D},\sigma^{2})}$ (15)
the acceptance ratio can be further simplified to,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime}|\mathcal{D},\sigma^{2})}{p(m|\mathcal{D},\sigma^{2})}}_{\begin{subarray}{c}\text{Ratio of model}\\ \text{posteriors}\end{subarray}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\Big{\}}$ (16)
This simplification is only available for particular choices of the prior $p(\Xi,m)$ that are conjugate, e.g. the Gaussian likelihood defined above (and defined independently to the prior over models). However, it gives direct access to the marginal posteriors $p(m|\mathcal{D})$ (see [23, 24] for example). Considering here a conjugate Gaussian prior over the parameter values (independently of their inclusion in the model), one has,
$p(\Xi|m)\triangleq\mathcal{N}(\mu_{m}^{(0)},\Sigma_{m}^{(0)})$ (17)
$\begin{split}p(m|\mathcal{D},\sigma^{2})&\propto p(m)p(\mathcal{D},\sigma^{2}|m,\Xi)\\ &\propto p(m)\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)\end{split}$ (18)
where,
$\Sigma_{m}=\sigma^{2}\left[\Theta_{m}^{\mkern-1.5mu\mathsf{T}}\Theta_{m}+\Sigma_{m}^{(0)-1}\right]^{-1}$ (19)
$\mu_{m}=\Sigma_{m}\Sigma_{m}^{(0)-1}\mu_{m}^{(0)}+\frac{1}{\sigma^{2}}\Sigma_{m}^{\mkern-1.5mu\mathsf{T}}\Theta^{\mkern-1.5mu\mathsf{T}}\dot{x}$ (20)
are the posterior mean and covariance of the parameters for a given model $m$ (see [23, 24] for example). Thus, the overall acceptance probability of a move from $\Xi$ to $\Xi^{\prime}$ can be written,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime})}{p(m)}}_{\text{Model prior}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{\sigma^{-k^{\prime}}|\Sigma_{m^{\prime}}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m^{\prime}}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m^{\prime}}\mu_{m^{\prime}}\right)}{\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)}}_{\begin{subarray}{c}\text{Model posterior ratio}\end{subarray}}\Big{\}}$ (21)
The above can now be readily used to sample from the required posterior. In practice, the noise variance in (3) will be unknown. This problem can be simply addressed by including a Gibbs move for the noise variance given the model and the parameters, which are available following a reversible-jump move. In this work, an inverse-gamma prior is used. The Gibbs step is therefore,
$p(\sigma^{2})\triangleq\mathcal{IG}(a^{(0)},b^{(0)})$ (22)
Then by the conjugacy of the prior and likelihood, the posterior over the noise variance is available exactly as an inverse-gamma distribution,
$p(\sigma^{2}|\Xi,m,\mathcal{D})=\mathcal{IG}(a,b)$ (23)
where,
$a=a^{(0)}+\frac{n}{2}$ (24)
$b=b^{(0)}+\frac{\sum_{i}^{n}(\dot{x}-\Theta(x)\Xi)^{2}}{2}$ (25)
where $n$ is the number of data in $\dot{x}$. The overall sampler is summarised in Algorithm 1.
Algorithm 1 Gibbs sampler with reversible-jump model moves
Parameter prior mean and variance $\mu_{m}^{(0)},\Sigma_{m}^{(0)}$, variance prior parameters $a^{(0)},b^{(0)}$, prior over models $p(m)$, jump transition probabilities $J(m^{\prime}|m)$.
Initialise the chain at $m=\mathcal{M}$ (i.e. with every term included) and $\sigma_{0}^{2}$
for Number of samples $N$ do
     Propose a model move $m\rightarrow m^{\prime}$ by sampling from $J(m^{\prime}|m)$
     Accept the model move with probability $\alpha(\Xi\rightarrow\Xi^{\prime})$ given by (21) else remain in place.
     Propose a new noise variance $\sigma^{2}$ using the Gibbs update given by (23)
end for
return samples from $p(\Xi,m|\mathcal{D})$
3.4 Related work
Several authors have approached the joint model structure and parameterisation problem from a Bayesian perspective. A key contribution is the work of Abdessalem et al. [8, 9] wherein an approximate Bayesian computation (ABC) method was applied to the identification of a nonlinear system from within a small number of candidate models. Although the results of the method are compelling (approximate posteriors, jointly over models and parameters), the utility of the method is limited by the extreme computational cost of ABC.
The proposed method is also not the first time that the SINDy method has been formulated in a Bayesian context. Previous attempts have made use of sparsity inducing priors [5, 25, 26, 27, 17, 28]. These priors are placed over parameters in $\Xi$ such that much of the posterior mass is concentrated at or around zero (thus removing them from the model with high probability). Although sparsity-inducing priors are shown to be very effective in correctly identifying model terms (by setting superfluous parameters to zero), the method cannot obtain a distributional estimate over the terms in the model independently of their parameterisation [^3] Because sparsity inducing priors are usually defined continuously over the parameter range it is not possible to marginalise out the parameter values without setting a heuristic threshold on which parameter values should be considered insignificant.. In contrast, the method proposed in this work differs in that prior distributions are placed over both model terms and their parameters independently. Another challenge with sparsity-inducing priors is that the inference is usually intractable and can be expensive to compute. The work in this paper is able to significantly reduce the computational burden by marginalising the parameters $\Xi$ from the acceptance ratio (after [22]) such that a new parameter need only be sampled when updating the noise variance.
A related work is the recent contribution of Fung et al. [18]. The authors propose a Bayesian formulation of the SINDy method by tracing uncertainty through the nonlinear library and the differential operator in order to account for noise on observed $x$. The authors then make Gaussian approximations to recover the model evidence in a Bayesian manner. The authors then propose to greedily select terms that maximally increase the model evidence until no more such terms are available. This results in a single model structure estimate with a posterior distribution over its parameterisation. The method proposed in this differs in that a joint posterior distribution over both model terms and their parameterisation is available explicitly, thanks to the reversible-jump sampler. The proposed approach is also not limited to greedily maximising the model evidence, which can be important in situations whereby the posterior is multimodal in model space.
The method presented herein is also not the first use of RJMCMC in the Bayesian identification of dynamic models. Troughton and Godsill [22] and Dahlin et al.  [29] consider an RJMCMC scheme for the identification of linear autoregressive models, and a RJMCMC scheme is employed in [30] to select between two candidate nonlinear models. However, to the author’s best knowledge, this work represents the first application of RJMCMC to the SINDy-type formulation of [1], and the first to explicitly enumerate a full posterior distribution over both model terms and their parameterisation jointly.
Another related contribution is the uncertainty quantification enabled by the ensemble-SINDy (E-SINDy) method of Fasel et al. [16]. E-SINDy enables heuristic UQ through bootstrapping and (robust) averaging (b(r)agging) on subsets of model terms and data. In this way, the authors are able to give heuristic estimates of quantities such as term inclusion probabilities and bootstrapped distributional estimates of model parameters. While these quantities are undoubtably useful to practitioners, their heuristic nature makes their theoretical interpretation difficult. Nevertheless, E-SINDy is a natural benchmark comparison for the method proposed in this work. In all forthcoming case studies, both methods are demonstrated.
4 Results
In order to demonstrate the effectiveness of the proposed method, three benchmark case-studies are presented here. Results from both the proposed method and E-SINDy [16] are presented and compared in terms of their confidence in identifying the true dynamics.
In all case-study examples, the prior structure of the inference is set up to be weakly uninformative. In particular, the values,
$\mu_{m}\triangleq\bm{0}^{(0)},\quad\Sigma_{m}^{(0)}\triangleq 10^{3}\times I$ (26)
$a^{(0)}\triangleq 0,\quad b^{(0)}\triangleq 0$ (27)
are chosen. The prior distribution over models is selected initially to be flat with $p(m)\propto 1\ \forall\ m\in\mathcal{M}$. The authors would note that this choice not promote model parsimony a priori. For the jump transitions $J(m^{\prime}|m)$ the authors propose a ‘bit-flipping’ scheme whereby in each proposal step a random index in $[1,n]$ (where $n$ is the number of rows in $\Theta$) is selected with equal probability. If that term is present in the current model then it is removed, if not then it is added. If a move is rejected then a new index is drawn, again with even probability. The chain is initialised with all terms in $\Theta$ present in the model and with the noise variance set to $\sigma^{2}=1$. The Gibbs sampler is run for $6\times 10^{3}$ steps in total, with the first thousand samples discarded to remove the effects of any transient ‘burn-in’ behavior. The authors remark that this length of chain appears more than adequate in terms of convergence to the target distribution.
The E-SINDy results presented hereafter are generated using the publicly available ‘pysindy’ library [19]. In all cases, $5\times 10^{3}$ models are sampled (in line with the length of the chains in the RJMCMC) and both data and library ensembling are activated, with the number of candidates dropped in each sample set to 1. In both cases, the sequentially-thresholded least-squares (STLSQ) algorithm is used to perform the sparse regression. For more detail on the E-SINDy method and for interpretation of these hyperparameters the interested reader is directed to the original article [16] and the ‘pysindy’ documentation [19].
4.1 Legendre polynomials
Figure 1: Comparison between RJMCMC and SINDy in quantifying model uncertainty for the toy polynomial data. In the plot, the bar-charts depict model inclusion probability (left axis), point plots depict samples of model parameters (right, log axis). Horizontal bars depict median parameter values, black diamonds depict true values of parameters in the underlying data-generating model.
Figure 2: Samples from the identified posterior distribution (BINDy, green) and ensemble (E-SINDy, orange) in both interpolation and extrapolation regimes.
The first case-study is a static sparse polynomial regression problem in which $\Theta(x)$ is static and is set to be the first 10 Legendre polynomials on the interval $x\in[-1,1]$. In order to demonstrate the recovery of small-valued parameters, a coefficient vector is selected with both zero and near-zero elements. The coefficient vector used here is given by,
$\Xi^{*}\triangleq\left[0.549,0,0.603,0.545,0.424,0.006,0,0,0,0.004\right]^{\mkern-1.5mu\mathsf{T}}$ (28)
The target of the regression is then generated as,
$y\triangleq\Theta(x)\Xi^{*}+\epsilon\quad\epsilon\sim\mathcal{N}(\bm{0},\sigma^{2}_{n}I)$ (29)
whereby Gaussian noise is added to the target data with standard deviation equal to 5% of the root-mean-square (RMS) of the nose-free data.
The performance of both E-SINDy and the proposed method are compared in Figure 1. As can be seen in the figure, Both the proposed method and E-SINDy methods are able to correctly assign low inclusion probability to the terms in $\Xi^{*}$ equal to zero. However, the STLSQ algorithm is unable to identify terms with parameters lower than the threshold parameter $\lambda$, which in this case study is set to the default value of 0.1. While this is a simplified example, it serves to demonstrate a particular pathology, which may appear when using parameter value as a proxy for term importance. However, the E-SINDy method shows multi-modality, producing some samples in clusters away from the true value. This can be attributed to the library bagging operation. Every subset of the library functions produces a sample (none are rejected - there is no Metropolisation step). This has the effect that some samples are generated with important terms missing, biasing the parameter values of the remaining terms.
4.2 Lynx-hare population dynamics
Figure 3: Comparison between RJMCMC and SINDy in quantifying model uncertainty for the lynx-hare population example. In the plot, the bar-charts depict model inclusion probability (left axis), point plots depict samples of model parameters (right, log axis). Horizontal bars depict median parameter values, black diamonds depict maximum likelihood values of parameters in the Lokta-Volterra dynamics.
Figure 4: Samples from the identified posterior distribution (BINDy, green) and ensemble (E-SINDy, orange) in both interpolation and extrapolation regimes for the lynx-hare population example.
The second case-study considers the well-studied lynx-hare population dynamics dataset. This dataset consists of the number of pelts collected for hares and lynxes annually between 1900 and 1920 by the Hudson Bay Company and is considered to reflect the population level of the two species [16]. As such, the dataset is often used as a canonical example of data generated by the Lokta-Volterra predator-prey population model. Let $L$ and $H$ be the number of Lynx and Hare in the population respectively, then their population dynamics are governed by the first order ordinary differential equation,
$\begin{split}\frac{dL}{dt}&=aL+bHL\\ \frac{dH}{dt}&=cH+dHL\end{split}$ (30)
Where $a,b,c,d$ are unknown constants in the model. The dataset represents a significant challenge in that only 21 measurements are available.
Here, the derivatives of the lynx and hare populations are computed numerically in accordance with the approach taken in [16], the columns of $\Theta$ are also normalised and a threshold value of 0.19 for $\lambda$ is used. For $\Theta(x)$, library of polynomial terms, including interactions up to third order is used. The authors note that this constitutes a more challenging identification task [^4] The sizes of the respective model spaces increase from $2^{6}$ to $2^{10}$ when considering terms up to third order. than the one presented in [16], where only terms up to second order are considered.
The distributions identified by the two methods on the lynx-hare dataset are depicted in Figure 3. In the figure, it can be seen that both models are assigning high probability to terms in the Lokta-Volterra model. However, the probability mass is much more concentrated around the these terms in the BINDy posterior, compared to those in the E-SINDy ensemble. Once again, it can be seen that the parameter values are tightly distributed around the maximum likelihood values for both methods (assuming Lokta-Volterra dynamics), although the E-SINDy method produces a number of samples close to the cutoff threshold. Particularly concerning, is that the median parameter value for all terms in the hare evolution (as identified by ensemble SINDy) is zero. This corresponds to the STLSQ algorithm in E-SINDy removing all parameters and returning a zero model much of the time. These constant dynamics (with no terms in the model) identified by the SINDy ensemble are evident in Figure 4 which plots time histories (integrated forward in time numerically, from known initial conditions) of samples from both BINDy and E-SINDy. By comparison, the samples from the BINDy posterior represent the dynamics well, given the scarcity of data available. In extrapolation, neither model is able to closely follow the mean dynamics (computed as the least squares estimate of $a,b,c,d$, assuming the model structure to be known). However, the uncertainty in the dynamics offered by the BINDy posterior is visually far more compelling, expanding in both the location and magnitude of the population peaks for each species–as might be expected. Access to meaningful uncertainty quantification for real datasets in the low-data regime is of huge importance, for example when modelling epidemiological dynamics, wherein predicting the height and timing of peaks is critical.
4.3 Lorenz attractor dynamics
Figure 5: Comparison between RJMCMC and SINDy in quantifying model uncertainty for the Lorenz attractor example. In the plot, the bar-charts depict model inclusion probability (left axis), point plots depict samples of model parameters (right, log axis). Horizontal bars depict median parameter values, black diamonds depict true values of parameters in the underlying data-generating model.
Figure 6: Samples from the identified posterior distribution (BINDy, green) and ensemble (E-SINDy, orange) in both interpolation and extrapolation regimes for the Lorenz attractor example.
The third case study considers the chaotic dynamics of the Lorenz attractor. The dynamics of the Lorenz attractor are well studied in the SINDy literature [1, 16], and have become a benchmark system for advancements to the SINDy method. The dynamics are driven by the following system of ordinary differential equations.
$\begin{split}\frac{dx}{dt}&=10(x_{2}-x_{1})\\ \frac{dy}{dt}&=x_{1}(28-x_{3})\\ \frac{dz}{dt}&=x_{1}x_{2}-\frac{8}{3}x_{3}\end{split}$ (31)
For the case study in this work, the Lorenz equations above are simulated for 10 seconds (with a further 5 seconds of unseen data for extrapolation) from an initial condition of $[x_{1},x_{2},x_{3}]^{\mkern-1.5mu\mathsf{T}}=[-8,7,27]^{\mkern-1.5mu\mathsf{T}}$ at a sample rate of $10^{2}$ Hz. Overall, $10^{3}$ points are available for training. As before, $\Theta(x)$ is set as a polynomial library, considering terms up to third order. Before training, noise at 2.5% RMS is added to the state variables. The state derivatives are computed numerically using a polynomial smoothed finite-difference scheme with a window size of 5 and a polynomial order of 3. For the SINDy results, the threshold parameter is set to 0.2 (as is done in [16]). All other parameters are unchanged from the previous case-studies.
For this third case study, the effect of the prior over models $p(m)$ is also considered. In addition to the flat prior used in the previous case-study examples, a regularising geometric prior over the total number of non-zero elements in the model is considered. The probability mass function of the regularising prior is given by,
$p(m)=(1-\theta)^{d}\theta$ (32)
where $d$ is the number of terms in model $m$, and $\theta$ is a hyperparameter, set to the strongly regularising value of 0.99 in this work. The effect of this prior over models is to strongly penalise (by assigning low prior probability) models with many terms.
The result of the UQ for the Lorenz equations is compared in Figure 5 for BINDy (with and without the regularising prior) and E-SINDy. It can be seen in the figure that the unregularised BINDy assigns high posterior probability to many spurious terms. In contrast, Both E-SINDy and the regularised BINDy assign high confidence to terms in the Lorenz equations. This is not an unexpected result. The nature of the Lorenz equations, and the use of the polynomial library results in many, highly-correlated basis functions in $\Theta(x)$. This, plus the effect of noise permit many incorrect (but low error) parametrisations of the observed dynamics. Both E-SINDy and regularised BINDy are able to overcome this difficulty by explicitly promoting model parsimony. Because the unregularised BINDy enforces no such cognitive bias, it is unsurprising that the identified posterior contains many terms. Summary statistics of the three identified posteriors are provided in Table 1.
As before, sampled model trajectories (integrated forward in time) are plotted for all three methods in Figure 6. The chaotic nature of the Lorenz equations makes tracking the true dynamics highly challenging. The well-known ‘lobe-switching’ behavior of the dynamics means that small errors can quickly cause significant deviations as the dynamics become governed by a different local attractor. This effect is clearly present in the identified distributions, which quickly become multimodal after a few oscillations. Qualitatively, it is observed that the regularised BINDy samples remain in step with the true dynamics for the longest, with E-SINDy the first to diverge. This divergence can be seen at $t<1$s where orage lines deviate from the observed data. Indeed, this is supported by the numerical results in Table 2 which show lower error statistics for the regularised BINDy approach. In the extrapolation region, the distributions have become very diffuse, in keeping with what might be expected for forecasting a chaotic attractor such as the Lorenz equations.
Table 1: Posterior results for the Lorenz attractor case-study. Term inclusion probabilities $p(\Theta_{i}(x)|\mathcal{D})$ and expected parameter values $E[\Xi_{i}|\mathcal{D}]$ (and standard deviations) for each of the three considered methods.
Equation Term BINDy E-SINDy BINDy (regularised)
$p(\Theta_{i}|\cdot)$ mean (std) $p(\Theta_{i}|\cdot)$ mean (std) $p(\Theta_{i}|\cdot)$ mean (std)
$10(x_{2}-x_{1})$ $x_{1}$ 1.0 -10.4 (0.754) 0.947 -8.76 (3.88) 1.0 -10.1 (0.0785)
$x_{2}$ 1.0 8.6 (0.594) 0.949 9.38 (2.28) 1.0 10 (0.0668)
$x_{1}(28-x_{3})$ $x_{1}$ 1.0 27.3 (0.473) 0.95 23.9 (9.32) 1.0 27.4 (0.31)
$x_{2}$ 0.998 -0.781 (0.229) 0.851 1.22 (5.79) 1.0 -0.845 (0.107)
$x_{1}x_{3}$ 1.0 -0.981 (0.0173) 0.948 -0.888 (0.266) 1.0 -0.983 (0.0073)
$x_{1}x_{2}-\frac{8}{3}x_{3}$ $x_{3}$ 0.998 -2.9 (0.577) 0.922 -2.48 (0.912) 1.0 -2.66 (0.0238)
$x_{1}x_{2}$ 1.0 1.03 (0.052) 0.931 0.936 (0.262) 1.0 0.996 (0.0039)
Table 2: Median, Mean and standard deviations of sample mean-squared errors computed from the trajectories in Figure 5.
Equation BINDy E-SINDy BINDy (regularised)
median mean std median mean std median mean std
$10(x_{2}-x_{1})$ 79.6 81.9 20.7 86.1 86.5 19.6 74.5 74.4 17.4
$x_{1}(28-x_{3})$ 108 111 25.7 114 117 27.3 101 99 21.8
$x_{1}x_{2}-\frac{8}{3}x_{3}$ 69.4 77.5 27 85.4 102 110 60.6 66.7 22.8
5 Discussion
In the first two case studies, strongly parsimonious models were identified by the proposed methodology. Counterintuitively, these models were assigned high posterior probability despite a flat prior over the model space. It is interesting to consider how this can be the case, for which the authors offer two explanations. On examination of the acceptance ratio in (21), one finds that the ratio is proportional to $\sigma^{-k^{\prime}}/\sigma^{-k}$. Intuitively, this ratio has a slight regularising effect that penalises larger models that have the same fit to the data. Another important consideration is the correlations between the columns of $\Theta(x)$ and the state derivatives $\dot{x}$. These correlations depend on the data, the system under study and chosen library of basis functions. Many, highly correlated terms can complicate the identification procedure with many models able to produce predictions close to the observed data. In contrast, in the case where only a few terms have high correlation, the identification of models with few terms is simplified. This effect could also explain the parsimonious models in the first two case studies.
As with any Bayesian approach, the specification of prior distributions plays an important role. A limitation of the current approach is that the priors over the parameters themselves are required to be conjugate. For most applications this is not expected to be an issue, although there are several situations (for example bounded parameter spaces or monotonic functions) where this requirement could become restrictive. The authors cannot presently envisage a solution to this issue that does not come at considerable computational cost, should alternative, non-conjugate priors be desired.
As has been demonstrated in the third case-study of this work, the choice of prior over models $p(m)$ may be critical in enforcing parsimony. Here, a strongly regularising geometric distribution parameterised by the number of terms in the model has been applied successfully. However, vastly many more parametrisations are possible. It is expected that prior selection in practice will be driven by expert domain knowledge, wherever possible. For example, one could set the prior over each term directly, assigning high prior probability to terms that are expected to be in the model and low probabilities otherwise. Alternatively one could promote parsimony in some types of library functions while permitting dense models in other types. The flexible nature of the Bayesian formulation presented in this work, allows modellers to encode both domain knowledge and cognitive biases a priori.
One could even imagine a library tempering scheme (similar in some ways to the sequential library and parameter bagging in [16]) whereby model terms with low probability are tempered out of the prior during sampling. Practically this would enable more efficient sampling (as more model moves are likely to be accepted) as well as permit the operator to start the sampler with a vastly greater (or even infinite) library of terms, with the expectation that all but those in the typical set would be tempered out. Such a scheme would bear some resemblance to well-studied sequential Monte-Carlo (SMC) samplers [31], for which a great deal of analysis has already been conducted. However, the application of SMC samplers to domains whereby parameters are permitted to change (such as in the interacting scheme of Jasra et al. [32]) has not been as widely studied and remains an open challenge.
A related area for future investigation is the choice of the ‘jump’ kernel $J(m^{\prime}|m)$. The ‘bit-flipping’ method applied herein has proved to be effective in the case-studies considered so far, but it might be ineffective when significant multi-modality is expected in the posterior, for example when very differently parameterised models describe the data well, low probability intermediate steps might inhibit exploration of the posterior. Instead, it is interesting to imagine more sophisticated schemes that could better suit problems of this type. Also interesting are so-called ‘non-reversible’ jump methods [33] that use an adaptive scheme to ensure that the posterior is explored efficiently.
A key assumption in SINDy methods is that the dynamics are described as $\dot{x}\approx\Theta(x)\Xi$. Implicit in the least-squares formulation of [1] (as well as almost every other approach for learning sparse $\Xi$) is the Gaussian noise model in (3). The assumption of Gaussian noise on $\dot{x}$ only, with $x$ (or at least $\Theta(x)$) noise free, is unlikely to hold for many practical applications – especially in the low-data, high-noise regime. The calculation of $\dot{x}$ numerically from noisy $x$ is likely to only compound this issue, particularly for second-order dynamical systems where the numerical derivative must be computed twice. It is certainly possible to envisage more advanced statistical descriptions of the noise model, whereby noise enters the observed $x$ is transformed nonlinearly by $\Theta$ and by the numerical differentiation in computing $\dot{x}$ (such an attempt is made in [18]). However, such a formulation is likely to make the inference described in this work impractically expensive. Despite the strong assumption of the noise model in 3, the authors are encouraged by the strong performance of the proposed method in both numerical and experimental case-studies. One potential explanation for this performance is that the noise variance parameter $\sigma^{2}$ is freely able to inflate during the sampling. This can have the effect of ‘swallowing’ additional variance that may be present in a more accurate likelihood model. The result is a Gaussian approximation to the likelihood function (in a manner similar to [18]).
6 Conclusion
In this work, a novel method for Bayesian identification of nonlinear dynamic systems in a SINDy-like manner has been proposed, considering a joint inference over both model terms and their parameterisation. It is the argument of this paper that such a posterior over models is more useful than single (or even an ensemble of) sparse models as it permits the operator to make probability-informed choices, and propagate meaningful uncertainty to further analyses.
Overall the proposed methodology has several advantages in heuristic quantification of uncertainty over ensemble-based SINDy, including prior choice over models and a valid distributional formulation. The Bayesian formulation has been shown to be effective in three benchmark case studies, correctly assigning high posterior probability to terms in the underlying differential equations.
Authors’ contributions
MDC, Conceptualisation (Equal), Methodology (Equal), Software (Lead), Data Curation (Lead), Writing - Original Draft (Lead), Writing - Review & Editing (Supporting)
Conflict of interest declaration
The authors declare there are no competing interests in this work.
Funding
The authors of this paper gratefully acknowledge the support of the UK Engineering and Physical Sciences Research Council (EPSRC) via grant references EP/W005816/1 and EP/W002140/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.